OpenStudy (julburg22):

Write 8pr/12p in simplest form (Will post actual fraction shape in comments.)

OpenStudy (flamingpiano):

I believe you can cancel out the p in the numerator and denominator.

OpenStudy (julburg22):

\[\left(\begin{matrix}8pr \\ 12p\end{matrix}\right)\]

OpenStudy (flamingpiano):

Then condense the #s

OpenStudy (julburg22):

I am thinking the answer is \[\left(\begin{matrix}2r \\ 3\end{matrix}\right)\]

OpenStudy (flamingpiano):

Divide the numbers by 4, giving you 2r/3, yes!

OpenStudy (julburg22):

Thank you!

OpenStudy (flamingpiano):

No problem. :)